Miguel invested his savings in two investment funds. The amount he invested in fund A was 3x as much as the amount he invested in fund B. Fund A returned a 7% profit and fund B returned a 2% profit. How much did he invest in fund B, if the total profit from two funds together was $2070?
Answer by mananth(16946)

Miguel invested his savings in two investment funds.
The amount he invested in fund A was 3x as much as the amount he invested in fund B......................x


Fund A returned a 7% = 7%(3x)
profit and fund B returned a 2% 2%(x)profit.
profit from two funds together was $2070
2%x+7%(3x)=2070
multiply by 100
2x+7x=207000
9x=207000
/9
x=23000
3*23,000 = 69,000
$23,000 at 2%
$69,000 at 7%